OpenStudy (anonymous):

The drawing shows two displacements, 3.0 km due north and 4.0 km due east, and the corresponding resultant vector R. What is the angle θ that R makes with respect to due north?

hero (hero):

vector R = 5.0 km tan x = 4/3 x = tan^-1(4/3) x = 53.13 (approx)